# Retrieval-Enhanced Knowledge Editing for Multi-Hop Question Answering in Language Models

### v1

@@ Abstract.
Large Language Models (LLMs) have shown proficiency in question-answering tasks but often struggle to integrate real-time knowledge updates, leading to potentially outdated or inaccurate responses. This problem becomes even more challenging when dealing with multi-hop questions since they require LLMs to update and integrate multiple knowledge pieces relevant to the questions. To tackle the problem, we propose the Retrieval-Augmented model Editing (RAE) framework tailored for multi-hop question answering. RAE first retrieves edited facts and then refines the language model through in-context learning. Specifically, our retrieval approach, based on mutual information maximization, leverages the reasoning abilities of LLMs to identify chain facts that naïve similarity-based searches might miss. Additionally, our framework incorporates a pruning strategy to eliminate redundant information from the retrieved facts, which enhances the editing accuracy and mitigates the hallucination problem. Our framework is supported by theoretical justification for its fact retrieval efficacy. Finally, comprehensive evaluation across various LLMs validates RAE’s ability in providing accurate answers with updated knowledge.
Model editing, question answering, retrieval-augmented generation
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Large language models (LLMs) exhibit impressive performance in question-answering (QA) tasks (Radford et al., 2019; Wang and Komatsuzaki, 2021; Touvron et al., 2023b), yet they face a significant challenge: their static knowledge base cannot be easily updated in real-time, leading to the risk of generating outdated or incorrect responses. To overcome this issue, model editing has been proposed for pre-trained LLMs to align their output with up-to-date knowledge (Cohen et al., 2023). Previous editing methods have shown to be effective in updating answers to single-hop questions (Meng et al., 2022a, b; Dai et al., 2021). However, it remains a challenging task for model editing to handle multi-hop questions, which is crucial to evaluate machine’s comprehension and reasoning abilities.
Answering multi-hop questions requires the integration of multiple pieces of knowledge. For instance, to answer the question ”What is the nationality of the author of ‘Harry Potter’?”, we must link two pieces of knowledge: ”(Harry Potter, author, J. K. Rowling)” and ”(J. K. Rowling, citizen of, United Kingdom)”, which collectively constitute a fact chain (Zhong et al., 2023). If we conduct a counterfactual edit on the first fact, i.e., “J. K. Rowling” is replaced by “Stephen King”, the subsequent knowledge must be adjusted accordingly, resulting in a totally different fact chain: ”(Harry Potter, author, Stephen King), (Stephen King, citizen of, United States).” Here, we use counterfactual edits to simulate real-world updates. Successful model editing for multi-hop questions requires that the edited LLMs identify and adopt the updated knowledge to derive the final answers.
@@
Figure 1. An example of the traditional similarity-based search that fails to retrieve the correct facts for LLM editing.
Unfortunately, existing editing methods are often limited in handling multi-hop questions. First, methods that alter model parameters, including fine-tuning (Zhu et al., 2020), locate-then-edit (Meng et al., 2022a, b), and meta-learning (Mitchell et al., 2021), suffer from the catastrophic forgetting issue (Dai et al., 2021; Zhai et al., 2023; Cohen et al., 2023), where the previously encoded knowledge could be lost after editing. Second, methods that depend on training auxiliary models also fall short in these scenarios. The auxiliary models are usually smaller language models, which lack the necessary reasoning capability to infer correct answers (Mitchell et al., 2022). In contrast, a third category of methods, based on Retrieval-Augmented Generation (RAG), modifies model outputs in a more effective manner (Zheng et al., 2023; Zhong et al., 2023; Cohen et al., 2023). These methods integrate updated knowledge directly into the model prompt, guiding LLMs through in-context learning (Zheng et al., 2023). RAG-based approaches have shown significant advantages since they are immune to the catastrophic forgetting problem, and the editing process can be conducted on the fly (Wang et al., 2023a; Han et al., 2023). Yet, the application of RAG to model editing with multi-hop questions still remains largely underexplored.
Furthermore, in practice, model editing often involves handling a large volume of edits, which are stored as fact triplets in a memory bank. In such cases, RAG-based multi-hop editing is expected to retrieve the most relevant facts for each question. However, this task is challenging for two reasons. First, extracting multi-hop facts requires the retriever to understand the complex connections among multi-relations within the question. A naïve design of the retriever is applying similarity-based search (Lewis et al., 2020; Wang et al., 2023a; Han et al., 2023) to obtain the top-$K$ facts that are most semantically similar to the question. However, semantic similarity alone does not guarantee that these facts contain the necessary information to correctly answer the question. We illustrate this issue with an example in Figure 1. The edited fact that should be retrieved is “[2]”: (Stephen King, citizen of, Canada), not “[4]”: (Harry Potter, citizen of, United States). The latter fact “[4]” is retrieved because it contains “Harry Potter” and “citizen of” that resemble our question “What is the nationality of Harry Potter’s author?”. Despite its higher similarity score, the fact “[4]” is actually irrelevant to the question. Therefore, effective retrieval requires a deep understanding of the question, a capability that goes beyond existing similarity-based search methods. Second, the retrieved knowledge could contain redundant information and degrade model editing performance. It is usually impractical to determine the exact amount of information to answer a specific question, so existing retrieval methods tend to return an extensive number of facts (Das et al., 2019; Yasunaga et al., 2021; Nie et al., 2020; Sidiropoulos et al., 2021) for more comprehensive coverage. While these methods do retrieve relevant facts, they also introduce redundant information. It is worth noting that incorporating irrelevant knowledge into the LLM input could mislead models and cause serious “hallucination” problems (Li et al., 2024, 2023), where LLMs will generate factually incorrect content based on noise instead of relevant facts. Thus, it is crucial to reduce noise before applying retrieved facts to prompts.
To bridge the gap, we propose a novel Retrieval-Augmented model Editing (RAE) framework, where we first retrieve edited facts and then refine the target model with these facts through in-context learning. To address the first challenge, we propose Mutual-information (MI) Maximization for edited facts retrieval. Here, MI quantifies the shared information between the target question and edited facts. Edited facts with higher MI are more relevant. However, directly computing MI is challenging. Thus, we decompose MI into a series of conditional probabilities, and we utilize the next-word prediction capabilities of pre-trained LLMs to estimate these probabilities. In this process, we use the target LLM’s contextual understanding and reasoning abilities to identify necessary facts for successful model editing. To tackle the second challenge, we propose Uncertainty-based Redundant Facts Pruning by utilizing LLM output confidence. Specifically, it selectively retains facts that increase the LLMs’ confidence in answering edited questions and discards irrelevant information. Finally, we theoretically justify the formulation of our retrieval objective. Overall, our main contributions are listed below:
  • We introduce a novel fact retrieval approach for multi-hop questions in model editing. This approach effectively harnesses the reasoning capabilities of LLMs to retrieve the most relevant multi-hop facts for each question.
@@
  • We propose a knowledge pruning strategy to reduce noise after the initial retrieval, mitigating the hallucination problem. Additionally, we provide theoretical analysis to justify our design for the retrieval objective.
  • We conduct extensive experiments on various language models in different sizes to verify the effectiveness of our proposed editing method. Empirical results demonstrate the superiority of our RAE framework compared to state-of-art baselines.
@@ 2.1. Model Editing for Single-hop Questions
In LLMs, a single model edit refers to updating a specific piece of factual knowledge (Mitchell et al., 2021; Meng et al., 2022b; Zheng et al., 2023). Each knowledge is defined as a triplet $\delta:=(h,r,t)$, where $h$, $r$, and $t$ denote the head entity, the relation, and the tail entity, respectively, such as (Harry Potter, author, J. K. Rowling). An edit is defined as changing the tail entity $t$ to a new entity $t^{\prime}$, i.e., $\delta\rightarrow\delta^{\prime}:=(h,r,t)\rightarrow(h,r,t^{\prime})$, where $\delta^{\prime}$ is the edited knowledge. Let $q$ denote the language model’s input. The goal of model editing is to modify a target model $f_{\theta}$, so that the new model $f_{\theta}^{\prime}$ produces an output $f_{\theta}^{\prime}(q)$ that is aligned with the new fact $\delta^{\prime}$, where $f_{\theta}^{\prime}(q)\neq f_{\theta}(q)$. Specifically, given $q=[h;r]$, the model is expected to output $t^{\prime}=f_{\theta}^{\prime}([h;r])$ if $h,r\in\delta^{\prime}$, where [;][;][ ; ] is the concatenation operator. However, if the input question is not relevant to the edit, i.e., $h\notin\delta^{\prime}$ or $r\notin\delta^{\prime}$, the model should output $t=f_{\theta}^{\prime}([h;r])$ that reflects the knowledge of LLMs before editing.
Table 1. Answering a 3-hop question $q$ with model editing. The pre-edited and edited answer are $t_{5}$ and $t_{3}^{*}$, respectively. $t_{5}$ and $t_{3}^{*}$ are the tail entity of $\delta_{5}$ and $\delta^{\prime}_{3}$. $G_{q}$ and $G^{*}_{q}$ denote the pre-edited and edited fact chain.
@@
$(\delta_{1}\rightarrow\delta_{1}^{\prime})$ Harry Potter, author, J. K. Rowling $\rightarrow$ Stephen King#edit
$(\delta_{2})$ Stephen King, citizen of, United States
$(\delta_{3}\rightarrow\delta_{3}^{\prime})$ United States, capital, Washington, D.C. $\rightarrow$ Boston    #edit
$(\delta_{4})$ J. K. Rowling, citizen of, United Kingdom
$(\delta_{5})$ United Kingdom, capital, London
@@
$(q)$ Which city is the capital of the country where the author of Harry Potter held citizenship?
@@
$(t_{5})$ London
$(t^{\prime}_{3})$ Boston
@@
$(\delta_{1})$ (Harry Potter, author, J. K. Rowling)
$(\delta_{4})$ (J. K. Rowling, citizen of, United Kingdom)
$(\delta_{5})$ (United Kingdom, capital, London)
@@
$(\delta_{1}^{\prime})$ (Harry Potter, author, Stephen King)     #edited fact
$(\delta_{2})$ (Stephen King, citizen of, United States)     #unedited fact
$(\delta_{3}^{\prime})$ (United States, capital,Boston)     #edited fact
2.2. Model Editing for Multi-hop Questions
Answering multi-hop questions presents a greater challenge. A multi-hop question seeks to identify a specific tail entity $t_{k}$ based on a sequence of linked facts: $\{(h_{1},r_{1},t_{1}),(h_{2},r_{2},t_{2}),...,(h_{k},r_{k},t_{k})\}$, where each tail entity is the head entity of the next fact: $t_{i}=h_{i+1}$. Each input question $q$ is associated with a sequence fact chain $G_{q}$ for question answering. A $k$-hop question is usually formulated using only the initial head entity $h_{1}$, and a series of relationships $\{r_{1},r_{2},...,r_{k}\}$. An example of model editing for a 3-hop question is shown in Table 1. Different fact chains are used to answer the question before and after editing. One key observation is that fact chains represent connected knowledge graphs, where a single entity is involved in two consecutive facts. Additionally, we notice a ”ripple effect” in these chains: An edit in the first fact $\delta_{1}$ will lead to changes in the subsequent facts, forming a new chain $G^{*}_{q}$. In practical scenarios, editing is usually conducted in batches, involving multiple fact changes simultaneously, resulting in an edited fact bank, denoted as $\Delta=\{\delta_{1}^{\prime},\delta_{2}^{\prime},...,\delta_{N}^{\prime}\}$, where $N$ is typically a large number (Meng et al., 2022b). Locating the relevant edited facts for one specific question is non-trivial due to the ”ripple effect”. To correctly answer multi-hop questions in model editing, it is crucial to address the retrieval task formally defined as follows:
@@ 2.2. Model Editing for Multi-hop Questions
Given an edited fact bank $\Delta=\{\delta_{1}^{\prime},\delta_{2}^{\prime},...,\delta_{N}^{\prime}\}$ with $N$ instances and a multi-hop question $q$ whose answer requires model editing, we want to retrieve its corresponding edited facts $\Delta_{q}=\{\delta_{i}^{\prime},\delta_{j}^{\prime},...,\delta_{k}^{\prime}\}$. The goal is to ensure the retrieved facts contain all the edited facts that appear in fact chain $G^{*}_{q}$, i.e., $\Delta_{q}\subseteq G^{*}_{q}$ and $\Delta\backslash\Delta_{q}\not\subset G^{*}_{q}$. Then, these facts are used to refine the target model $f_{\theta}$ to conduct model editing.
@@ 3. Methodology
Our Retrieval-Augmented Editing (RAE) framework, as shown in Figure 2, contains two key steps: (1) retrieving edited facts relevant to the target question, and (2) editing the language model using these retrieved facts via in-context learning. We will first discuss step (2) with the motivation of our design in the following section. The details of step (1) can be found in Sections 3.2 and 3.3.

### v2

# Retrieval-enhanced Knowledge Editing in Language Models for Multi-Hop Question Answering

@@ Abstract.
Large Language Models (LLMs) have shown proficiency in question-answering tasks but often struggle to integrate real-time knowledge, leading to potentially outdated or inaccurate responses. This problem becomes even more challenging when dealing with multi-hop questions, since they require LLMs to update and integrate multiple knowledge pieces relevant to the questions. To tackle the problem, we propose the Retrieval-Augmented model Editing (RAE) framework for multi-hop question answering. RAE first retrieves edited facts and then refines the language model through in-context learning. Specifically, our retrieval approach, based on mutual information maximization, leverages the reasoning abilities of LLMs to identify chain facts that traditional similarity-based searches might miss. In addition, our framework includes a pruning strategy to eliminate redundant information from the retrieved facts, which enhances the editing accuracy and mitigates the hallucination problem. Our framework is supported by theoretical justification for its fact retrieval efficacy. Finally, comprehensive evaluation across various LLMs validates RAE’s ability in providing accurate answers with updated knowledge. Our code is available at: https://github.com/sycny/RAE.
Model editing; question answering; retrieval-augmented generation
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management; October 21–25, 2024; Boise, ID, USA.booktitle: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management (CIKM ’24), October 21–25, 2024, Boise, ID, USAisbn: 979-8-4007-0436-9/24/10doi: 10.1145/3627673.3679722ccs: Information systems Question answeringccs: Computing methodologies Knowledge representation and reasoning
@@ 1. Introduction
Figure 1. An example of traditional similarity-based search that fails to retrieve the correct facts for LLM editing.
Large language models (LLMs) excel in answering factual questions due to their pre-training on extensive corpora (Radford et al., 2019; Wang and Komatsuzaki, 2021; Touvron et al., 2023b). However, their dependence on pre-trained knowledge can lead to outdated answers, which requires updates through model editing techniques (Meng et al., 2022a; Yao et al., 2023; Zhang et al., 2024a; Zhong et al., 2024). Among them, editing for multi-hop questions is particularly important, as real-world questions often require combining multiple knowledge pieces. For example, to answer ”Who is married to the British Prime Minister?”, one must connect several related facts, creating a fact chain, like ”(United Kingdom, head of government, Theresa May), ”(Theresa May, spouse, Philip May)”. If we update ”Theresa May” to ”Rishi Sunak” to reflect real-world change (wik, [n. d.]), it requires adjusting the related facts accordingly, resulting in a new fact chain: ”(United Kingdom, head of government, Rishi Sunak), (Rishi Sunak, spouse, Akshata Murty)”. Recent research shows that retrieval-augmented generation (RAG) is effective in LLM editing for multi-hop question answering (Zhong et al., 2023; Wang et al., 2024), surpassing other methods such as fine-tuning (Zhu et al., 2020) and locate-and-edit (Meng et al., 2022a, b) approaches. The RAG-based methods first retrieve relevant facts related to the question and then feed these facts into LLMs through in-context learning. Studies have shown that LLMs are highly receptive to external knowledge, even when it contradicts their internal pre-trained knowledge (Xie et al., 2023; Kortukov et al., 2024). Therefore, RAG can effectively update knowledge within LLMs but also avoid problems such as catastrophic forgetting (Gu et al., 2024a; Gupta et al., 2024) and hallucinations (Gekhman et al., 2024). However, simply using multi-hop questions as queries in RAG often fails to retrieve pertinent facts due to the complexity of the questions involved, as illustrated in Figure 1.
To address this, some methods (Wang et al., 2024; Zhong et al., 2023) propose breaking down complex multi-hop questions into simpler, single-hop queries to improve the effectiveness of the retrieval process. Despite enhancements, critical issues persist: 1) Commonly used models with fewer parameters (e.g., Llama2-7B (Touvron et al., 2023a)) exhibit significantly worse editing performance (more than 30% gap) compared to more powerful models like GPT-3.5 or GPT-4 (Zhong et al., 2023); 2) The performance of existing editing methods degrades as the number of editing cases increases, making them impractical for large-scale editing tasks (Zhong et al., 2023; Wang et al., 2024). The possible reason behind their performance degradation is that their editing effectiveness heavily relies on the quality of sub-questions generated for querying. Generating accurate sub-questions is challenging for models with less strong reasoning and planning capabilities. Inaccurate sub-questions will result in irrelevant fact retrieval, which may mislead the LLMs and reduce editing effectiveness (Yoran et al., 2023).
Since successful multi-hop editing depends on accurate retrieval of the question-specific facts, instead of applying the existing ”generate then retrieve” approach, we propose to directly fetch the required facts from the database by leveraging the next-token prediction capabilities of LLMs. Below are the details of our editing method.
1. The connection between edited facts helps retrieval: Each edited fact can be represented as a triplet of ”(head entity, relation, tail entity)”. The nearby facts in a fact chain are connected through entities: the tail entity of one fact becomes the head entity of the next. This observation inspires us to adopt a knowledge graph (KG) for storing these triplets. In a KG, each entity has only a limited number of neighboring entities, which narrows down the retrieval choices to a feasible number, rather than having to search through a vast database. Our method differs from existing approaches that store edited facts as embeddings in a vector database, whose editing performance decreases as the number of edits grows (Zhong et al., 2023; Wang et al., 2024).
2. Next-token prediction helps next-fact retrieval: LLMs are inherently skilled at predicting the next token in a sequence. In our study, we extend this capability to predict the next fact in a fact chain. Our approach first feeds LLMs a sequence that includes the question, any preceding fact, and the relevant entity. Then, the model predicts this entity’s next possible logical relation within the context of its input question. Finally, we retrieve the tail entity from the KG, based on the head entity and relation, to complete the fact chain. However, directly predicting the most probable next fact can lead to biases toward frequently occurring words (Welbl et al., 2021; Sun et al., 2019). Therefore, we predict facts that share the most relevant information with the question and preceding facts, using mutual information (MI) as our retrieval metric. We develop a technique to break down the MI into conditional probabilities that LLMs can effectively approximate (Radford et al., 2019), thus improving the accuracy of our predictions.
3. LLM internal state helps reduce redundancy: Irrelevant facts to the editing will mislead LLMs. Thus, we propose to prune redundant facts extracted by the retrieval step using the LLM’s prediction uncertainty. The uncertainty is minimized when the LLM is prompted with a correct fact chain, and increases when provided with incomplete or excessive facts. We quantify this uncertainty using the LLM’s output entropy. Unlike traditional methods that either limit the number of retrieval attempts or simply ask the model to identify redundant facts (Zhong et al., 2023; Wang et al., 2024), our approach provides a more effective way to ensure accuracy in the editing process.
@@
Overall, we name our approach as Retrieval-Augmented model Editing (RAE), where we introduce a novel fact retrieval method for multi-hop questions in model editing. We also propose a knowledge-pruning strategy to reduce noise after the initial retrieval, mitigating the hallucination problem. Additionally, we provide theoretical analysis to justify our design for the retrieval objective.
@@ 2.1. Model Editing for Single-hop Questions
In LLMs, a single model edit refers to updating a specific piece of factual knowledge (Mitchell et al., 2021; Meng et al., 2022b; Zheng et al., 2023). Each knowledge is defined as a triplet $\delta:=(h,r,t)$, where $h$, $r$, and $t$ denote the head entity, the relation, and the tail entity, respectively, such as (Misery, author, Stephen King ). An edit is defined as changing the tail entity $t$ to a new entity $t^{\prime}$, i.e., $\delta\rightarrow\delta^{\prime}:=(h,r,t)\rightarrow(h,r,t^{\prime})$, where $\delta^{\prime}$ is the edited knowledge. Let $q$ denote the language model’s input. The goal of model editing is to modify a target model $f_{\theta}$, so that the new model $f_{\theta}^{\prime}$ produces an output $f_{\theta}^{\prime}(q)$ that follows the new fact $\delta^{\prime}$, where $f_{\theta}^{\prime}(q)\neq f_{\theta}(q)$. Specifically, given $q=[h;r]$, $h,r\in\delta^{\prime}$, the model is expected to output $t^{\prime}=f_{\theta}^{\prime}([h;r])$, where [;][;][ ; ] is the concatenation operator. However, if the input question is not relevant to the edit, i.e., $h\notin\delta^{\prime}$ or $r\notin\delta^{\prime}$, the model should output $t=f_{\theta}^{\prime}([h;r])$ that reflects the original knowledge of LLMs.
2.2. Model Editing for Multi-hop Questions
Answering multi-hop questions presents a greater challenge. A multi-hop question seeks to identify a specific tail entity $t_{k}$ based on a sequence of linked facts: $\{(h_{1},r_{1},t_{1}),(h_{2},r_{2},t_{2}),...,(h_{k},r_{k},t_{k})\}$, where each tail entity is the head entity of the next fact: $t_{i}=h_{i+1}$. Answering each input question $q$ requires a fact chain $G_{q}$. A $k$-hop question can be formulated using only the initial head entity $h_{1}$, and a series of relationships $\{r_{1},r_{2},...,r_{k}\}$. An example of model editing for a 3-hop question is shown in Table 1. Here, we use counterfactual edits to simulate real-world updates. Different fact chains are used to answer the question before and after editing. One key observation is that fact chains represent connected knowledge graphs, where a single entity is involved in two consecutive facts. Additionally, we notice a ”ripple effect” in these chains: An edit in the first fact $\delta_{1}$ will lead to changes in the subsequent facts, forming a new chain $G^{*}_{q}$. In practice, knowledge editing is usually conducted in batches, involving multiple fact changes simultaneously, resulting in an edited fact bank $\Delta=\{\delta_{1}^{\prime},\delta_{2}^{\prime},...,\delta_{N}^{\prime}\}$, where $N$ is a large number (Meng et al., 2022b). Locating the relevant edited facts for a question is non-trivial due to the ”ripple effect”. To correctly answer multi-hop questions after model editing, it is crucial to address the retrieval problem formally defined as follows:
Table 1. Answering a 3-hop question $q$ with counterfactual model editing. The pre-edited and edited answer are $t_{5}$ and $t_{3}^{*}$, respectively. $t_{5}$ and $t_{3}^{*}$ are the tail entity of $\delta_{5}$ and $\delta^{\prime}_{3}$. $G_{q}$ and $G^{*}_{q}$ denote the pre-edited and edited fact chain.
@@ 2.2. Model Editing for Multi-hop Questions
$(\delta_{1}\rightarrow\delta_{1}^{\prime})$ Misery, author, Stephen King $\rightarrow$ Richard Dawkins#edit
$(\delta_{2})$ Richard Dawkins, citizen of, United Kingdom
$(\delta_{3}\rightarrow\delta_{3}^{\prime})$ United Kingdom, capital, London $\rightarrow$ Birmingham    #edit
$(\delta_{4})$ Stephen King, citizen of, United States
$(\delta_{5})$ United States, capital, Washington, D.C.
@@
$(q)$ Which city is the capital of the country where the author of Misery held citizenship?
@@
$(t_{5})$ Washington, D.C.
$(t^{\prime}_{3})$ Birmingham
@@
$(\delta_{1})$ (Misery, author, Stephen King)
$(\delta_{4})$ (Stephen King, citizen of, United States)
$(\delta_{5})$ (United States, capital, Washington, D.C.)
@@
$(\delta_{1}^{\prime})$ (Misery, author, Richard Dawkins)     #edited fact
$(\delta_{2})$ (Richard Dawkins, citizen of, United Kingdom)     #unedited fact
$(\delta_{3}^{\prime})$ (United Kingdom, capital,Birmingham)     #edited fact
@@
Given an edited fact bank $\Delta=\{\delta_{1}^{\prime},\delta_{2}^{\prime},...,\delta_{N}^{\prime}\}$ with $N$ instances and a multi-hop question $q$ whose answer requires model editing, we want to retrieve its corresponding edited facts $\Delta_{q}=\{\delta_{i}^{\prime},\delta_{j}^{\prime},...,\delta_{k}^{\prime}\}$. The goal is to ensure that all the edited facts necessary for answering $q$ are retrieved, i.e., $\Delta_{q}\subseteq G^{*}_{q}$ and $\Delta\backslash\Delta_{q}\not\subset G^{*}_{q}$. Then, these facts are used to refine the target model $f_{\theta}$ for editing.
@@ 3. Methodology
Our Retrieval-Augmented Editing (RAE) framework, as shown in Figure 2, contains two key steps: (1) retrieving edited facts relevant to the question, and (2) editing the language model using these retrieved facts via in-context learning. We will first discuss step (2) with the motivation of our design in the following section. The details of step (1) are in Sections 3.2 and 3.3.